1.A=P+PRT SOLVE FOR P
2.A TOY ROCKET IS LAUNCHED UPWARD WITH AN INITIAL VELOCITY OF 96 ft/sec. AFTER t sec ITS VELOCITY IS v IS 96-32t ft/sec. a) SOLVE FOR t IN TERMS OF v. b) HOW LONG AFTER LAUNCHING WILL THE ROCKET TAKE TO REACH ITS HIGHEST POINT, THAT IS, WHEN IS THE VELOCITY 0ft/sec?
3.AVERAGE OF THREE NUMBERS: A= a+b+c DIVIDED by 3. SOLVE FOR a
4.ARITHMATIC PROGRESSION: S=a+(n-1)d. SOLVE FOR a
5.dx-ey=f SOLVE FOR y WHEN d=5, X=-1 e=-2, AND f=7

CAN YOU PLEASE SHOW ME HOW TO SOLVE THESE?

Answer by stanbon(48551)

1. A=P+PRT SOLVE FOR P
A=P(1+RT)
P=A/(1+RT)
----------------
2.A TOY ROCKET IS LAUNCHED UPWARD WITH AN INITIAL VELOCITY OF 96 ft/sec. AFTER t sec ITS VELOCITY IS v IS 96-32t ft/sec. a) SOLVE FOR t IN TERMS OF v. b) HOW LONG AFTER LAUNCHING WILL THE ROCKET TAKE TO REACH ITS HIGHEST POINT, THAT IS, WHEN IS THE VELOCITY 0ft/sec?
V=96-32t
Add 32t to both sides to get:
32t+V=96
Subtract V from both sides to get:
32t=96-V
Divide by 32 to get:
t=(96-V)/32
--------------
3.AVERAGE OF THREE NUMBERS: A= (a+b+c) DIVIDED by 3. SOLVE FOR a
A = (a+b+c)/3
Multiply both sides by 3 to get:
3A=a+b+c
a=3A-b-c
---------------
4.ARITHMETIC PROGRESSION: S=a+(n-1)d. SOLVE FOR a
a=S-(n-1)d
---------------
5.dx-ey=f SOLVE FOR y WHEN d=5, X=-1 e=-2, AND f=7
y=(dx-f)/e
y=(5(-1)-7)/(-2)
y=(-12)/(-2)
y=6
-----------
Cheers,
Stan H.